OpenStudy (anonymous):

Could someone please help me? I'm trying to figure out the problem, but none of the answer choices are what my final answer is coming out to be. Find the angle between the given vectors to the nearest tenth of a degree. u = <2, -4>, v = <3, -8> A. 3.0 Degrees B. 6.0 Degrees C. -7.0 Degrees D. 16.0 Degrees I keep getting around 0.99

jimthompson5910 (jim_thompson5910):

I'm assuming you computed \[\Large \frac{u \cdot v}{|u|*|v|} = 0.9945054529214\] correct?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

Apply the arccosine function to both sides to isolate theta \[\Large \cos(\theta) = \frac{u \cdot v}{|u|*|v|}\] \[\Large \cos(\theta) = 0.9945054529214\] \[\Large \arccos(\cos(\theta)) = \arccos(0.9945054529214)\] \[\Large \theta = ???\]

OpenStudy (anonymous):

Oh! So the answer would be 6??

jimthompson5910 (jim_thompson5910):

yeah arccos(0.9945054529214) = 6.00900595749781